# Retrieval-Based Video Language Model for Efficient Long Video Question Answering

### v1

@@ Appendix B More visualization results
We visualize more examples from the QAEgo4D and WildQA datasets in Fig. 3 and Fig. 4, with the following information. 1) The first row shows the video chunk samples by uniformly selecting 5 video chunks. 2) The second row shows the retrieved 5 chunks (ordered by time order) in our R-VLM. We mark the groudtruth chunks by red box. 3) We show the learned similarity score curve based on which the top $K$ chunks are selected. The horizontal axis represents the identity of chunk and the vertical axis denotes the similarity score of that chunk. The groundtruth chunks and our retrieved chunks are also marked. 4) The question and answers from different models: R-VLM, R-VLM w/Uni., Video-ChatGPT, and Video-LLaMA, respectively.
@@
We also show some failure examples in Fig. 5. A detailed analysis of the reasons for failure is given in the figure caption. There are two main cases of failure. One is that the retrieval does not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer is wrong. For the later cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
Appendix C Computational Complexity
@@
The computational cost comes from two parts. The first part is to encode the video frames through the CLIP encoder and the spatial-temporal pooling to get chunks. The second part is the retrieval of K=5 chunks and put them to LLM for inference. The spatial-temporal pooing and retrieval is very fast and negligible. On a single A100, we tested 120 60s videos from Social-IQ 2.0 and calculated the average inference time cost for a video. For a single video, the first part for vision feature extraction takes an average of 0.14s (in parallel for 60 frames), and the second part takes an average of 2.42s. The total time is 2.56s. Actually, for an even longer video, the time consumption for the second part does not increase since the input number of vision tokens is fixed (i.e., 68$\times$5=340) in our scheme, which is favored for long video or streaming video understanding. The GPU memory consumption is about 17GB. Note that the computational cost for the LLM is proportional to the number of tokens.
The FLOPs for LLM inference can be roughly estimated as 2PD, where P denotes the number of parameters (model size), and D denotes the number of tokens. The computational complexity of LLM is proportional to the number of tokens which consists of text tokens (question and answer) and vision tokens. The LLM model size P is 6.7B. On the training dataset, the average number of tokens for question and answers is 80, i.e., $D_{tex}=80$. This varies on different datasets. For simplicity, we assume the number is the same for all the datasets. We denote the number of vision tokens as $D_{vis}$. The total number of tokens is $D=D_{tex}+D_{vis}$. For the four video datasets, WildQA, QaEgo4D, lifeQA, Social-IQ 2.0, the average number of vision chunks is 19, 122, 20, and 16, where each chunk has 68 tokens. Thanks to the retrieval, only $K=5$ chunks ($D_{vis}^{\prime}=5\times 68=340$ tokens) instead of all the chunks are needed as the input to LLM. Therefore, the computational cost (FLOPs) for LLM inference can be saved approximately $\frac{D_{vis}-D_{vis}^{\prime}}{D_{tex}+D_{vis}}$, which are 69% (i.e., $(19\times 68-5\times 68)/(80+19\times 68)$), 95%, 71%, and 64%, respectively.
(a) We can see that the gray car does not appear in the uniformly sampled video chunks. Our R-VLM correctly answers that the car was parked in the parking lot (outdoors), but R-VLM w/Uni.’s answer was the garage (indoors). Video-LLaMA does not answer where the car is and the groundtruth frames do not appear in the frame 100. Video-ChatGPT made the similar mistake as R-VLM w/Uni.
@@
(b) The broom is small and is on the left in the red boxed image. Our R-VLM captures exactly where the broom is, i.e., on the ground. R-VLM w/Uni. does not capture the video chunks with broom and thus does not answer accurately. The answer of Video-ChatGPT is irrelevant to the question. The answer from Video-LLaMA is redundancy and tedious, where the mentioned blue jean jacket and boat actually do not appear in the video.
Figure 3: Visualization of video QA examples from QAEgo4D.
(a) The uniform sampling miss the chunks for sharpening process in GT-segs (at the beginning of video). As a result, LLM does not see the knife running along the leather, and only see the knife and some delicate small objects. Therefore, R-VLM w/Uni. mistakenly thought that this individual was carving patterns or making designs. Our retrieved chunks retain the process of the knife running on the leather and therefore R-VLM gives the correct answer. Both Video-LLaMA and Video-ChatGPT answered that people are cutting leather with a knife to make art, rather than sharpening the knife.
(b) In this video, collecting the kindling takes a short time, while placing the tinder on the stones takes a longer time. Uniform sampling makes LLM think that there is no process of collecting kindling and output the wrong answer of “digging a hole”. Our R-VLM identified the relevant chunks of “gathering” even though those chunks only take a small duration in the entire video, generating correct answer. Video-LLaMA’s prediction is not accurate since in fact the man did not gather deadwood and leaves in the video.
Figure 4: Visualization of video QA examples from WildQA.
@@
(a) A failure case from WildQA. This is a video of a person firing art. Although our method R-VLM retrieved the correct chunks, it gave the wrong answer of “cook food”. We think this is due to the visual ambiguity of the target object and the biases of the LLM.
@@
(b) A failure case from WildQA. Groundtruth chunks correspond to the chunks where three types of animals present, namely dolphins (the first chunk), seagulls, and human. Our method only retrieved the seagull and human chunks, but missed the dolphin chunk. R-VLM provided wrong answer due to the imperfect retrieval and the unsatisfactory reasoning capability of the used LLM.
@@
(c) A failure case from QAEgo4D. Our method did not find the correct chunks. Therefore, large language model did not correctly answer the question and provided hallucinated answer.
@@
alt="[LOGO]">

### v2

@@ IV-E Visualization Analysis
1) The first row shows the video chunk examples by uniformly selecting 5 video chunks. 2) The second row shows the retrieved 5 chunks (arranged in chronological order) in our R-VLM. We mark the groudtruth chunks by red box. 3) We show the curve of the learned similarity score (see Eq. (2)) based on which the top $K$ chunks are selected. The horizontal axis represents the indices of chunks and the vertical axis denotes the similarity score. The groundtruth chunks and our retrieved chunks are also marked. 4) The question and answers from different models.
We can see that the predicted answers of our R-VLM is more accurate than Video-ChatGPT, and Video-LLaMA. For QAEgo4D, the average duration is about 8 minutes (120 video chunks), whereas the groundtruth segments span only 2% of the entire video on average. It is difficult to hit groundtruth video segments by uniformly sampling $K(=5)$ chunks among 120 video chunks. Fig. 2 shows that the fragments selected by uniform sampling are in general irrelevant to the problem. Our learnable retrieval can accurately find the segments where the answer lies in. Feeding the correct chunks to the LLM makes it possible to obtain the correct answer to the question.
@@
Besides the successful cases, we present some failed cases in Fig. 9 of Supplementary Material B-E. There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the latter cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
IV-F Computational Complexity
TABLE XI: Average time cost for 60 seconds videos from Social-IQ 2.0. The first part is feature extraction and the second part is retrieval and LLM inference.
Feature Extraction Retrieval & LLM inference Total
Time (seconds) 0.14 2.42 2.56
The computational cost of our framework mainly comes from two parts. The first part is to encode the video frames through the CLIP encoder and the spatial-temporal pooling to get chunks. The second part is the retrieval of $K=5$ chunks and put them to the LLM for inference.
@@ IV-F Computational Complexity
On a single A100, we tested 120 videos with 60 seconds each from Social-IQ 2.0 and calculated the average inference time for a video, as shown in Table XI. For a video, the first part for vision feature extraction takes an average of 0.14 seconds (in parallel for 60 frames), and the second part (mainly LLM inference) takes an average of 2.42 seconds. The total time is 2.56 seconds. The retrieval by encoding the question representation, and calculating the similarity scores between the question representation and the chunks, is very fast with almost negligible time. Note that for a longer video, the time consumption for LLM inference does not increase since the input number of visual tokens to the LLM is fixed (i.e., 68$\times$5=340) in our scheme, which is favored for long video or streaming video understanding.
TABLE XII: Analysis of FLOPs saving of our scheme for LLM inference. $Avg.chunks$ denotes the average number of video chunks. $S_{vis}$ denotes the average number of visual tokens of all the chunks of a video (obtained by multiplying $Avg.Chunks$ by the number of visual tokens per chunk (i.e., 68)). $S_{vis}^{\prime}$ denotes the average number of visual tokens of the selected $K=5$ chunks of a video (5 $\times$ 68 = 340).
WildQA QaEgo4D lifeQA Social-IQ 2.0
$Avg.~{}chunks$ 19 122 20 16
$S_{vis}$ 1292 8296 1360 1088
$S_{vis}^{\prime}$ 340 340 340 340
$Save$ 69% 95% 71% 64%
[…]
$FLOPs\_Saved=\frac{S_{vis}-S_{vis}^{\prime}}{S_{tex}+S_{vis}}.$ (5)
We show the computation saving of four long video datasets in Table 5. Thanks to our retrieval, only $K=5$ chunks ($S_{vis}^{\prime}=5\times 68=340$ tokens) instead of all the chunks are needed as the input to LLM. The saved computational cost (FLOPs) ranges from 69% to 95%, demonstrating the high efficiency of our solution.
Moreover, our retrieval-based solution can significantly save the memory consumption. For an A100 GPU of 80G memory, given extracted chunk features, when we set batch size to 40 and $K$ to 5, the memory occupation is about 70G in training. In contrast, when a baseline scheme without using retrieval that all the chunks (instead of 5 selected chunks) are input to LLM, videos longer than 40 seconds are not able to be filled to the memory (out of memory) in training. In contrast, our scheme could support very long videos (even 80+ hours).
V Conclusion
The comprehension of long videos using LLMs remains an under-explored area. There are two main challenges associated with comprehending long videos. 1) Long videos generally lead to abundant visual tokens, which increase computational cost for LLM inference. 2) Global aggregation of visual tokens inevitably results in the loss of vision details especially when the question relevant video chunks take only a small portion of the entire video. Moreover, question irrelevant chunks introduce interference. In this work, we address these issues by introducing a simple yet effective retrieval-based video-language model (R-VLM) for long-video understanding. Our R-VLM effectively reduces the number of video tokens, preserves the most informative information, eliminates noise interference, and thus enhances system performance. Our experimental results demonstrate the effectiveness of our designs for comprehending long videos.
This paper is the result of an open source research project starting from 2023.
Appendix A More Details on Visual Token Extraction of a Chunk
To better illustrate the procedure of visual token extraction of a chunk, Fig. 3 illustrates the spatial and temporal pooling for obtain 68 visual tokens for a chunk.
Appendix B More Experimental Results
B-A More Details on Evaluation Metrics
In this paper, we follow the metrics of accuracy and average score as proposed by Video-ChatGPT [1] for performance evaluation, where we use ChatGPT (gpt-3.5-turbo) to assist in judging the correctness of model predictions. ChatGPT accepts questions, groundtruth answers, and the model predictions as input. For each question-answer pair, ChatGPT gives a binary judgment of “yes” or “no” to identify whether the predicted answer is correct or not, for accuracy evaluation. Moreover, an integer score of 0-5 is also given by ChatGPT to indicate how similar the prediction is to the answer. 0 represents the lowest score and 5 represents the highest score.
B-B Human Evaluation Results
We have conducted human studies in 100 samples of four test datasets on two models, the baseline scheme Video-ChatGPT, and our scheme R-VLM. We asked three people to blindly rate correct/wrong and scores and show the average evaluation results in Table R1-1. We can see that our method R-VLM achieves higher accuracy and score than Video-ChatGPT under human evaluation. The main trends from ChatGPT evaluation are consistent with our human evaluation but the values and relative gains between two schemes are different, which also indicates ChatGPT evaluation has a certain degree of reliability but are not perfect.
TABLE XIII: The results of human evaluation and ChatGPT evaluation on 100 samples, which are randomly selected from four test datasets.
Evaluator Video-ChatGPT R-VLM (Ours)
ChatGPT 48/2.98 50/3.03
Human 42.33/2.26 46.33/2.41
B-C More Ablation Studies
Comparison with Q-Former-based Designs We compare our retrieval-based design with Q-former designs under our framework. 1) We use a Q-Former to globally aggregate all the chunk tokens to be the input of the LLM, where the input to Q-Former is the $L_{i}\times(N+M)$ tokens, where $L_{i}$ denotes the number of chunks, $N+M=68$ denotes the reduced number of tokens (through spatial-temporal pooling) in a chunk. The output is 340 visual tokens. There is no retrieval in this scheme, and we dub it VLM-QFormer. Table XIV shows the comparison with our scheme. We can see that our R-VLM outperforms VLM-QFormer significantly, thanks to the retrieval-based design which can exclude the inference of irrelevant visual tokens. 2) Similar to InstructBLIP [21], we use instruction-aware Q-Former to globally aggregate all the chunk tokens to be the input of the LLM. The difference with VLM-QFormer is that the instruction is also taken as the input to Q-Former to learn instruction (question)-aware visual tokens. We dub this scheme as VLM-InstructQFormer. We can see that VLM-InstructQFormer outperforms VLM-QFormer on most datasets but is inferior to our R-VLM, demonstrating the effectiveness of our retrieval-based design for extracting question-relevant visual features.
TABLE XIV: Comparison with Q-Former based two designs under our framework, evaluated in terms of accuracy (%)/score.
VLM-Qformer VLM-InstructQformer R-VLM (Ours)
ActivityNet-QA 49.34/3.35 48.94/3.33 49.65/3.34
WildQA 58.59/3.16 58.28/3.17 64.82/3.39
QaEgo4D 30.58/2.35 31.23/2.41 32.51/2.45
lifeQA 35.48/2.56 37.10/2.62 38.71/2.61
Social-IQ 2.0 54.30/3.14 57.62/3.19 63.65/3.40
Figure 3: Illustration of the spatial and temporal pooling to obtain 68 visual tokens for a chunk. We perform spatial average pooling with stride 2 to have $M$$\times$ $\bar{h}$$\times$ $\bar{w}$ = $4\times 8\times 8=256$ tokens per chunk, where $\bar{h}$=$h/2$ and $\bar{w}$=$w/2$. This is equivalent to taking the CLIP features of reduced resolution as the extracted feature. Global spatial average pooling for each frame is performed and thus we obtain $M$ tokens for the $M$ frames ($M=4$). Temporal pooling for each spatial position is performed to have $N=\bar{h}\times\bar{w}=8\times 8=64$ tokens. Therefore, we have $N+M=64+4=68$ tokens for a chunk.
B-D More Results on Other Datasets
We also report the results on the subset of EgoSchema in Table XV. It can be seen that our R-VLM method achieves the best performance.
TABLE XV: Performance comparison with the state-of-the-art methods for video QA on the EgoSchema subset of 500 testing pairs.
Model Acc.
ViperGPT [19] 15.8
Video-ChatGPT [1] 20.0
SeViLA [24] 25.7
mPLUG-Owl [22] 33.8
Video-LLaVA [44] 36.8
R-VLM (Ours) 46.0
B-E More Visualization Results
We visualize more examples from the QAEgo4D and WildQA datasets in Fig. 5, Fig. 5 with the following information.
1) The first row shows the video chunk examples by uniformly selecting 5 video chunks. 2) The second row shows the retrieved 5 chunks (arranged in chronological order) in our R-VLM. We mark the groudtruth chunks by red box. 3) We show the curve of the learned similarity score (see Eq. (2)) based on which the top $K$ chunks are selected. The horizontal axis represents the indices of chunks and the vertical axis denotes the similarity score. The groundtruth chunks and our retrieved chunks are also marked. 4) The question and answers from different models.
For the WildQA dataset, groundtruth segments usually locate at different locations in the video and last for a period of time. Although uniform sampling sometimes hits a certain groundtruth segments, there is no guarantee. In contrast, our learned retrieval can correctly hit the segments where the groundtruth segments are located on. In this way, the LLM can better understand the video and answer some detailed questions (refer to “vegetation types” in Fig. 5).
In addition to successful cases, we present some failed cases in Fig.9 of Appendix B-E. There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the latter cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
@@ B-E More Visualization Results
Figure 4: Visualization of a video QA example from QAEgo4D. The broom is small and is on the left in the red boxed image. Our R-VLM captures exactly where the broom is, i.e., on the ground. R-VLM w/Uni. does not capture the video chunks with broom and thus does not answer accurately. The answer of Video-ChatGPT is irrelevant to the question. The answer from Video-LLaMA is redundancy and tedious, where the mentioned blue jean jacket and boat actually do not appear in the video.
Figure 5: Visualization of video QA examples from WildQA. In this video, two clips show vegetation and the remaining clips show mountains, rivers, etc. Uniform sampling mainly obtains segments such as mountains and rivers rather than segments with vegetation. Therefore, only the terrain was answered, without giving vegetation types. In contrast, our retrieved chunks contain video clips of vegetation. Thus the types of vegetation are predicted correctly: trees, bushes, forest. Video-ChatGPT gives a global description and does not answer specific vegetation types.
@@
Figure 6: A failure case from WildQA. This is a video of a person firing art. Although our R-VLM retrieved the correct chunks, it gave the wrong answer of “cook food”. This maybe caused by the visual ambiguity of the target object and the biases of the LLM.
@@
Figure 7: A failure case from WildQA. Groundtruth chunks correspond to the chunks where three types of animals present, namely dolphins (the first chunk), seagulls, and human. Our method only retrieved the seagull and human chunks, but missed the dolphin chunk. R-VLM provided wrong answer due to the imperfect retrieval and the unsatisfactory reasoning capability of the used LLM.
@@
Figure 8: A failure case from QAEgo4D. Our method did not find the correct chunks. Therefore, large language model did not correctly answer the question and provided hallucinated answer.
@@
B-F More Discussion on Memory Cost
We have also estimated the memory cost for training different schemes when we set the batch size as 20 and use input video sequences of 80 seconds. Our scheme with $K$=5 uses similar size of memory (45.3 GB) as VideoChat-GPT (42.2 GB). In contrast, the total memory cost without using retrieval (input all the chunks, in total 1360 visual tokens) is 75.1GB, demonstrating that our retrieval-based design (45.3GB) can significantly save memory and is efficient.
B-G Discussion on the Effectiveness of Retrieval
There are no ground-truth locations of the question relevant chunks for supervision. The effectiveness of our learned chunk selection (retrieval) benefits from four aspects that promote the question (text) and vision alignment. First, the use of pretrained CLIP vision encoder and CLIP text encoder, where their features are trained to be matched/aligned. Second, the learnable MLP block (which follows the CLIP text encoder) makes it possible for the question and vision alignment, which the original CLIP pretraining (caption and vision alignment) has not been. As shown in Table VII in our main manuscript, R-VLM w/ CLIP M. denotes the use of the final CLIP class token feature of vision and text for matching instead of learnable retrieval. We can see that our final scheme R-VLM with learnable matching outperforms R-VLM w/ CLIP M. significantly. Third, the proposed soft matching (SM) loss promotes the alignment of question and vision features. R-VLM w/o SM denotes the scheme without using SM loss, which is inferior to R-VLM. Forth, the end-to-end instruction tuning (loss from the LLM) facilitates the selection of right chunks through the gradient back-propagation. Intuitively, only when the chunks are selected correctly, does the model generate the right answers. R-VLM w/o G. denotes our scheme but without back-propagating the gradient from the LLM loss. R-VLM that allows such back-propagation to the learnable MLP block outperforms R-VLM w/o G. obvious. We can see that all these factors jointly make the retrieval-based solution feasible and efficient.
Appendix C Limitation and Future Work
We have explored the retrieval-based solution for the long video understanding. We experimentally found that using $K=5$ selected chunks can generate superior performance. Sample adaptive $K$ should deliver optimal results but we leave this as future work.
When the retrieval does not select the correct video chunks, our model will fail to generate correct answers as some failed cases in Fig.9 (b) and (c). As a remedy, we could incorporate some global video tokens (e.g., like Video-ChatGPT) to assure the preservation of global information to alleviate this. We leave this as future work.
Appendix D Impact Statement
This paper presents work that aims to advance the field of video-language modeling. Our research is designed to be a positive force for innovation purpose. However, future applications should be mindful to avoid malicious uses. When we release code and models, we would set a safeguard by requiring that users adhere to usage guidelines.